OpenStudy (anonymous):

Find functions f and g so that h(x) = (f o g)(x) h(x) = 10/square root of 7x+6

OpenStudy (anonymous):

How do I solve this? Somebody please show me easy steps. I have an exam tomorrow and I need to understand!

OpenStudy (anonymous):

you have a lot of choices here

OpenStudy (anonymous):

i assume it is \[f\circ g(x)=\frac{10}{\sqrt{7x+6}}\]

OpenStudy (anonymous):

you could say \(g(x)=7x+6,f(x)=\frac{10}{\sqrt{x}}\) then \[f(g(x))=f(7x+6)=\frac{10}{\sqrt{7x+6}}\]

OpenStudy (anonymous):

you could say \[g(x)=\sqrt{7x+6},f(x)=\frac{10}{x}\] then \[f(g(x))=f(\sqrt{7x+6})=\frac{10}{\sqrt{7x+6}}\]

OpenStudy (anonymous):

you could say \[g(x)=\frac{1}{\sqrt{7x+6}},f(x)=10x\] then \[f(g(x))=f(\frac{1}{\sqrt{7x+6}})=\frac{10}{\sqrt{7x+6}}\]

OpenStudy (anonymous):

you can probably come up with more if you like but there is no one right answer,there are several ways to do it